an archer stands on a cliff elevated at 45 m high from the ground and shoots an arrow at the angle of 35 degrees above the horizontal with the speed of 100ms-1.how long does it fly in the air? how far from the base of the cliff does the arrow fly until it hits the ground? calculate the speed of the arrow just before it hits the ground.
1
Expert's answer
2021-08-16T08:36:19-0400
"0=45+100\sin{35}t-0.5(9.8)t^2\\t=12.4\ s"

"x=100(12.4)\cos{35}=1020\ m"

"v^2=100^2+2(9.8)(45)\\v=104\frac{m}{s}"